OpenStudy (anonymous):

solve for x in the given interval 2tanx=secx (-2pi,0)

OpenStudy (anonymous):

\(\large 2tanx=secx \) \(\large 2\frac{sinx}{cosx}=\frac{1}{cosx} \) \(\large \frac{2sinx}{cosx}=\frac{1}{cosx} \) \(\large 2sinx=1 \) numerators must be equal \(\large sinx=\frac{1}{2} \) Can you do the rest from here?

OpenStudy (anonymous):

Yess thank you :D